A solution is prepared by adding 30g of sugar into 500g of water.How
much sugar must be added to the solution to obtain a new solutionn
containing 40% of sugar?

Answer by richwmiller(17219)

.40(530+x)=0.056604*(530)+x
0.4x+212. = x+30
182.-0.6x = 0
182.=0.6x
x = 303.333